### v1

# MunchSonic: Tracking Fine-grained Dietary Actions through Active Acoustic Sensing on Eyeglasses

Abstract.
We introduce MunchSonic, an AI-powered active acoustic sensing system integrated into eyeglasses, designed to track fine-grained dietary actions such as hand-to-mouth movements for food intake, chewing, and drinking. MunchSonic emits inaudible ultrasonic waves from a commodity eyeglass frame. The reflected signals contain detailed information about the positions and movements of various body parts, including the mouth, jaw, arms, and hands, all involved in eating activities. These signals are processed by a custom deep learning pipeline to classify six actions: food intake, chewing, drinking, talking, face-hand touching, and other activities not belonging to the tracking set (null). In an unconstrained user study with 12 participants, MunchSonic achieved a 93.5% macro F1-score in a user-independent evaluation with a 2-second time resolution, demonstrating its effectiveness. Additionally, MunchSonic accurately tracks eating episodes and the frequency of food intake within those episodes.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 International Symposium on Wearable Computers; October 8–12, 2024; Melbourne, Australiabooktitle: Proceedings of the 2024 International Symposium on Wearable Computers (ISWC ’24), October 5–9, 2024, Melbourne, Australiaprice: 15.00doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/XX/XX
[…]
Research
1. Introduction
Tracking dietary activity is crucial for measuring nutrition levels and managing chronic diseases. However, the lack of reliable automated dietary monitoring systems forces most people to rely on manual input-based applications, which are prone to self-reporting errors. To address this, researchers have developed wearable dietary activity recognition systems. Despite advancements in sensing modalities and machine learning algorithms to process the data, tracking fine-grained dietary actions such as food intake, chewing, and drinking, while distinguishing them from similar body movements like talking, or face touching, as well as from non-eating activities, remains a significant challenge for the wearable computing research community.
Dietary activities such as eating and drinking require coordinated hand, jaw, and mouth movements. Detecting fine-grained dietary actions necessitates the simultaneous tracking of both hand and mouth movements. Current state-of-the-art dietary monitoring systems, whether on eyeglasses (Shin et al., 2022; Bedri et al., 2020; Mirtchouk et al., 2017), earables (Bi et al., 2018), or smartwatches (Zhang et al., 2022), require sensing modalities at multiple body locations to achieve this, which can be costly, power-inefficient, and cumbersome for many users.
Recently, the active acoustic sensing system ActSonic (Mahmud et al., 2024) showed promising efficacy in tracking everyday activities, including identifying eating and drinking episodes, using a single point of instrumentation on eyeglasses. However, this system has not yet been able to predict fine-grained dietary actions, such as chewing or hand-to-mouth intake gestures, which are related to important biomarkers for longitudinal health monitoring (Slyper, 2021), such as chewing patterns, drinking habits, or snacking behaviors. In this paper, we aim to address the following research question:
  • Can an active acoustic sensing system instrumented on smart glasses be used to further distinguish fine-grained dietary actions in unconstrained environments beyond only identifying eating moments?
To address this challenge, we developed MunchSonic, a fine-grained dietary action recognition system in the form factor of commodity eyeglasses using active acoustic sensing. Similar to ActSonic (Mahmud et al., 2024), our system employs two pairs of microphones and speakers attached to the hinges of the eyeglasses. MunchSonic transmits inaudible ultrasonic waves that reflect off various body parts, generating a two-dimensional echo profile, which serves as a range spectrogram of the scanned body parts. This profile is then analyzed by a lightweight deep learning framework to infer fine-grained dietary actions. Our system is based on the observation that fine-grained dietary actions involve movements of the mouth, jaw, and arms, which can be inferred from reflected ultrasound, as shown in prior work (Mahmud et al., 2023; Li et al., 2022, 2024). Therefore, we believe it is possible to distinguish these actions from the reflected signals beyond merely identifying eating or drinking moments (Mahmud et al., 2024). We evaluated our system in a user study with 12 participants in unconstrained environments of their choice. The results showed that MunchSonic can accurately recognize fine-grained dietary actions such as hand movements for intake, chewing, and drinking, as well as non-dietary actions like talking, face touching, and other activities (null), achieving an F1-score of 93.5%.
The contributions of the MunchSonic system are as follows:
  • Advanced the knowledge of wearable-based active acoustic sensing by demonstrating the feasibility of using active acoustic sensing on glasses to recognize fine-grained dietary actions beyond only identifying eating moments.
  • Evaluation of the proposed system through a 12-participant user study in unconstrained environments, with ground truth data annotated for each second.
2. Related Work
Year Study Wearable
Sensing
Modalitie(s)
Detected
Event(s)
Evaluation
Metric
Power
Consumption
2015 Thomaz et al. (Thomaz et al., 2015b)
Wrist-mounted
Device
Microphone
(Passive Audio)
Meals F1-score = 79.8% -
2016 Farooq et al. (Farooq and Sazonov, 2016a) Eyeglasses Piezoelectric
Food
Intake
F1-score = 100.0%* -
2016 iHear Food (Gao et al., 2016)
Bluetooth
Headset
Microphone
(Passive Audio)
Eating
Episodes
Accuracy = 76.82% -
2017 EarBit (Bedri et al., 2017)
Outer Ear
Interface
IMU, IR Proxomity
Sensor, Microphone
Chewing F1-score = 80.1% -
2017 GlasSense (Chung et al., 2017) Eyeglasses Load Cell
Head Movement,
Talking, Chewing
F1-score = 94.0%* -
2017 Mirtchouk et al. (Mirtchouk et al., 2017) Google Glass
IMU, Microphone,
Motion Sensor
Meals
Precision = 31%
Recall = 87%
2018 Auracle (Bi et al., 2018) Earpiece
Contact
Microphone
Eating
Episodes
F1-score = 77.5%
14.47 mW
2018 Zhang et al. (Zhang and Amft, 2017) Eyeglasses EMG Electrode
Eating
Event
F1-score = 77% 81.96 mW
2020 FitByte (Bedri et al., 2020) Eyeglasses
1 Camera,
1 Proximity
Sensor, 6 IMUs
Eating and
Drinking
Episodes
Precision = 82.8% (Eat-
-ing), 56.7% (Drinking)
Recall = 93.8% (Eating),
65.5% (Drinking)
105.08 mW
2022 MyDJ (Shin et al., 2022) Eyeglasses
Accelerometer,
Piezoelectric
Eating
Episodes
F1-score = 92% 26.06 mW
2024 MunchSonic Eyeglasses
Active Acoustic
Sensing
Intake, Chewing, Drink-
-ing, Talking, Face Touch
F1-score = 93.50% 96.5 mW
Table 1. Comparison of eating detection studies based on wearable form factors, sensing modalities, granularity of detected events, evaluation metrics (F1-score or precision/recall), and power draw (in mW). Asterisks (*) in the Evaluation Metric column indicate that the system was evaluated in lab studies; otherwise, the evaluation was conducted in free-living conditions.
Various wearable devices with different sensing modalities have been proposed to track eating events. This section first discusses non-eyeglass wearables, followed by eyeglasses. Table 1 summarizes prior studies on tracking eating activities.
Form Factors Other than Eyeglasses. Sensing systems for detecting eating behaviors have been integrated into various wearable devices. Smartwatch-based systems for eating (Thomaz et al., 2015a; Ye et al., 2016; Dong et al., 2013; Kyritsis et al., 2019; Sen et al., 2018; Zhang et al., 2022) and drinking (Hamatani et al., 2018), worn on the user’s dominant hand, mainly use Inertial Measurement Units (IMUs) to track hand-to-mouth gestures. These systems face challenges in distinguishing similar hand motions, leading to a high incidence of false positives in real-world applications (Tong et al., 2020; Chun et al., 2018a). Ear-worn devices, or earables, are also used to track eating events due to their proximity to the jaw, which moves during chewing. Techniques such as in-ear proximity sensing (Bedri et al., 2015a, b), passive acoustic sensing (Gao et al., 2016; Bi et al., 2018), or a combination of both (Bedri et al., 2017), have been employed to monitor chewing activity. Neckbands (Farooq et al., 2014; Rahman et al., 2014; Yatani and Truong, 2012) and necklaces (Alshurafa et al., 2015; Kalantarian et al., 2015; Chun et al., 2018b; Shin et al., 2019; Zhang et al., 2020b) use passive acoustic and proximity sensing to track sounds and motions associated with chewing and swallowing. Despite these advancements, necklaces, and neckbands have not gained the same popularity as other wearables (Amores et al., 2019).
Eyeglasses Form Factor. Dietary monitoring using eyeglasses is of particular interest because sensors located near the mouth and jaw can capture significant movement during eating and drinking. Additionally, eyeglasses have higher social acceptance (Kosmyna et al., 2019) (64% adoption among the US population (Farooq and Sazonov, 2018)) compared to other wearables. Researchers have attached Electromyography (EMG) electrodes to eyeglasses to monitor eating behaviors (Zhang and Amft, 2017). Piezoelectric sensors (Farooq and Sazonov, 2016a, b), which make skin contact, have been used to track jaw movements but the performance tends to be impacted under conditions of sweating or rapid user movement. Inertial sensors (Farooq and Sazonov, 2018) and load cells (Chung et al., 2017) have been proposed for dietary event tracking but have not shown promising results. Sensor fusion approaches (Bedri et al., 2020; Rahman et al., 2016; Mirtchouk et al., 2017; Bedri et al., 2022; Shin et al., 2022), integrating gyroscopes, accelerometers, proximity sensors, and microphones on eyeglasses, have shown performance improvements in tracking eating events. Passive acoustic sensing using contact microphones (Bi et al., 2018) captures chewing sounds but requires close contact between the sensor and the skin. Acoustic sensing-based activity recognition systems have included labels for eating (Mahmud et al., 2024) or drinking (Mollyn et al., 2022; Mahmud et al., 2024), but they focus on eating episodes rather than fine-grained actions.
In summary, compared to these prior work, MunchSonic is the first to infer fine-grained dietary actions by employing active acoustic sensing on eyeglasses to capture information related to fine-grained jaw and hand movements. This approach allows for detailed dietary action extraction, such as intake and chewing, unlike other systems summarized in Table 1, which primarily track the moments of eating activities.
3. System Implementation
The goal of MunchSonic is to capture the information related to movements on multiple body parts involved in dietary activities using an eyeglasses form factor. Previous research has shown active acoustic sensing’s efficacy in tracking facial muscle movements (Li et al., 2022, 2024; Zhang et al., 2023), upper body limbs (Mahmud et al., 2023), respiration (Wang et al., 2018), sign language gestures (Jin et al., 2021, 2023), and everyday activities (Mahmud et al., 2024). Inspired by these prior work, especially (Mahmud et al., 2024), MunchSonic integrates active acoustic sensing into eyeglasses to track these fine-grained dietary actions. This section discusses the active acoustic sensing mechanism, the hardware implementation, and the deep learning framework for processing the captured data.
3.1. Sensing Mechanism
The design of MunchSonic’s sensing system is inspired by the active acoustic sensing approach from (Mahmud et al., 2024). In essence, MunchSonic uses a similar active acoustic sensing processing method, based on Cross-correlation based Frequency Modulated Continuous Wave (C-FMCW) (Wang et al., 2018). These chirps transmitted from eyeglasses’ hinges, with frequency ranges of 18-21.5 KHz and 21.5-24.5 KHz for the left and right transmitters, respectively. The receiver samples at 50 KHz, with each chirp containing 600 samples and a sweep period of 0.012 seconds. The system can detect changes as small as 3.43 mm, given the slow speed of mouth, jaw, and hand movements compared to the speed of sound (343 m/s). The sensing range is up to 2.058 meters, though a shorter range is used for fine-grained dietary actions.
MunchSonic’s signal processing pipeline involves computing the cross-correlation of transmitted and received ultrasonic waves. The received signal is first filtered through a bandpass filter (18-21 KHz and 21.5-24.5 KHz) to eliminate audible frequencies. The cross-correlation output, called the Echo Profile, functions as a Range-FFT using acoustic C-FMCW waves. The Echo Profile is processed as a two-dimensional tensor ($x$-axis: time, $y$-axis: distance from MunchSonic eyeglasses form factor). To capture body part movements and eliminate static object reflections, the first derivative of the Echo Profile, termed the Differential Echo Profile, is used as input for the deep learning architecture to track fine-grained dietary actions.
3.2. Hardware and Form Factor
Figure 1. MunchSonic Hardware and Form Factor: (a) User wearing eyeglasses form factor, (b) Customized controller unit with nRF52840 microcontroller, (c) Top view of the eyeglasses form factor, (d) MunchSonic transceiver for active acoustic sensing housing one speaker (top) and one microphone (bottom).
We replicated the sensing system from ActSonic (Mahmud et al., 2024) with minor changes to the sensors and controller unit, using the OWR-05049T-38D speaker and ICS-43434 microphone (TDK, [n. d.]). Our customized controller unit, shown in Figure 1(b), features the nRF52840 microcontroller (Semiconductors, [n. d.]), two MAX98357A audio amplifiers, a BLE SGW1110 (SGWireless, [n. d.]) module, power management modules, and an SD card slot with a SanDisk Extreme Pro microSD card (Amazon.com, [n. d.]), optimized for power efficiency. The transceiver boards, each with a speaker and microphone, are attached to the eyeglass hinges, with placement refined for optimal tracking of mouth, jaw, and hand movements. The controller unit and LiPo battery are attached to one leg of the glasses frame, connected via Flexible Printed Circuit (FPC) cables and a JST connector. The system draws 96.5 mW of power, with a voltage of 4.02V and current of 24.0 mA, and the prototype used in the MunchSonic study, featuring a 290 mAh LiPo battery, lasts approximately 11.25 hours in SD card storage mode.
3.3. Deep Learning Framework
3.3.1. Data Processing and Model Architecture
We designed a lightweight deep learning framework to process the active acoustic data captured by the MunchSonic system. The differential echo profile described in 3.1 serves as the input to the deep learning model. We create overlapping sliding windows from the stream of differential echo profile data. To augment the data, we apply Gaussian noise to 5% of the sliding windows and mask 5% of the distance axis ($y$-axis) of the sliding window to simulate environmental noise and unrelated movements.
The preprocessed sliding windows are fed into a MobileNetV2 (Sandler et al., 2018) convolutional neural network encoder. This lightweight encoder, effective for edge devices, generates a 256-dimensional embedding vector for each sliding window. This vector is then fed into a classifier network with three feedforward layers (128, 64, and 6 neurons respectively). Each layer, except the last one, includes batch normalization (Ioffe and Szegedy, 2015), dropout (Srivastava et al., 2014) with a probability of 0.25, and Leaky ReLU (Xu et al., 2015) activation. Finally, a softmax operation provides the class probability distribution.
3.3.2. Model Training
The input to MunchSonic’s deep learning framework consists of sliding windows of differential echo profiles. The optimal sliding window for the MunchSonic dataset is 2 seconds long with 50% overlap. This window length was chosen because it appropriately fits an eating intake gesture. The optimal sensing range is 150 pixels on the $y$-axis of the differential echo profile, corresponding to 51.45 cm from the MunchSonic device. With speakers and microphones co-located on each hinge of the eyeglasses and transmitting acoustic waves in two frequency ranges, there are four channels in the differential echo profile (two direct and two indirect paths). Thus, the shape of the input sliding window is $(4\times 150\times 166)$, as each second of the sliding window contains $\lfloor\frac{50000}{600}\rfloor=83$ samples.
We use focal loss (Lin et al., 2017) for training the MunchSonic model due to its effectiveness in handling class imbalance in the MunchSonic dataset. The Adam (Kingma and Ba, 2014) optimizer with a cosine annealing learning rate scheduler is employed, starting with an initial learning rate of $10^{-2}$. The model, implemented using PyTorch and PyTorch Lightning frameworks, is trained for 30 epochs with a batch size of 128 on GeForce RTX 2080 Ti GPUs.
4. User Study
We conducted a 45-minute user study to evaluate MunchSonic in an unconstrained setting. This section details the study design and the resulting dataset.
4.1. Design of User Study
The active acoustic sensing-based activity recognition system ActSonic (Mahmud et al., 2024) has proven effective in tracking eating and drinking episodes in real-world settings over long-duration studies. For example, it can identify eating and drinking moments with a one-second resolution, achieving F1 scores of 90% and 88% respectively, without needing training data from new users in new environments. Building on this capability, the MunchSonic user study was designed with the assumption that active acoustic sensing on glasses can already accurately distinguish eating moments from other activities. This paper focuses on further extracting fine-grained dietary actions, such as hand-to-mouth movements for food intake, chewing, and drinking, as well as distinguishing eating from non-eating actions from active acoustic signals once eating moments are recognized. Specifically, the MunchSonic user study aimed to evaluate the system’s performance in unconstrained conditions in tracking fine-grained dietary actions and differentiating these from non-eating actions. We adopted a relatively shorter-duration free-living study, leveraging ActSonic’s proven efficacy in longer-duration in-the-wild studies. Additionally, ground truth annotation in our study was done per second, unlike other eating detection systems (Shin et al., 2022) that rely on self-reporting, enabling us to recognize dietary actions with a high time resolution of two seconds.
Figure 2. Foods consumed by user study participants.
4.2. Study Protocol
We conducted a user study approved by our organization’s Institutional Review Board (IRB) with 13 participants, 6 identifying as male and 7 as female, with an average age of $26.5$ years, ranging from 21 to 36 years. The participants wore the MunchSonic eyeglasses form factor and a chest-mounted GoPro HERO9 camera (GoPro, 2020) facing upward to capture their eating activities around the face as the ground truth. The camera recorded 720p video at 30 fps with a diagonal field of view of $148^{\circ}$. Note that one participant (P09) accidentally turned off the chest-mounted camera, resulting in no ground truth data for that participant. Therefore, we discarded that participant from the study and evaluated the system on the remaining 12 participants.
Participants were asked to bring a meal (breakfast, lunch, or dinner) from any source (home-cooked or restaurant) and come to the lab. After briefing them about the study procedure, we equipped them with the MunchSonic eyeglasses and the chest-mounted camera. We verified the data collection system with a short 5-minute session. Once verified, the data collection system was started, and participants were free to go anywhere while wearing the eyeglasses and camera. The only requirement was to finish the meal they brought to ensure sufficient samples of dietary movements. Apart from that, they could continue with their daily routine. The total duration of this unconstrained study was 45 minutes. Upon completion, participants returned to the lab to return the devices and were reimbursed with a \$20 USD gift card.
4.3. Dataset Statistics
We collected 540 minutes of data from 12 participants. Two researchers labeled the reference video from the chest-mounted camera at each second using the ANU-CVML Video Annotation Tool (Vidat) (Zhang et al., 2020a). The data was categorized into six classes: hand-to-mouth movement for food intake, chewing, drinking, and talking, face touch, and a null class for other activities. A synchronization script was used to extrapolate ground truth labels for the differential echo profile sliding windows. The class distribution was 53.5% null, 5.9% hand-to-mouth for food intake, 22.0% chewing, 2.2% drinking, 15.1% talking, and 1.3% touching face with hand for non-eating activities.
5. Performance Evaluation
We computed precision, recall, and macro F1-score to assess MunchSonic’s ability to track fine-grained dietary actions. Evaluations were conducted at two levels: frame-level inference for each two-second sliding window and episode-level inference for intake count and chewing time estimation. Employing a leave-one-participant-out evaluation strategy enabled user-independent assessment, using data from one participant as the test set and the remaining 11 for training and validation. This evaluation demonstrates the system’s performance without requiring training data from new users, which makes the system easier to deploy at scale.
5.1. Evaluation of Frame-Level Inference
Figure 3. Precision and recall of leave-one-participant-out evaluation of MunchSonic, where data from each participant on the $x$-axis serves as the test set.
We summarize the results of the leave-one-participant-out evaluation of sliding window prediction in Figure 3. According to the evaluation, the average macro F1-score across all 12 participants is 0.935 with a standard deviation of 0.031. It is evident that the MunchSonic system can track fine-grained dietary events with precision and recall of more than 90% for most participants. P08 demonstrates the worst performance in terms of both precision and recall. Additionally, P04 and P13 yielded slightly lower recall compared to the average. Our analysis of the ground truth videos for those participants suggests that the slightly worse performance can be attributed to the participants touching their faces several times during the study for activities not related to eating. This phenomenon is also evident in the confusion matrix in Figure 4. The performance degradation of MunchSonic for face-touch actions can be related to the imbalanced nature of the dataset collected in a completely uncontrained manner, which contains very few samples of this particular activity. Nonetheless, MunchSonic exhibits robust performance across all activities, with a mean precision and recall of $0.941\pm 0.027$ and $0.930\pm 0.041$ respectively, across all 12 participants.
Figure 4. Normalized confusion matrix of the leave-one-participant-out evaluation across 12 participants of the user study of MunchSonic.
5.2. Evaluation of Episode-Level Inference
Figure 5. Episode-level evaluation of MunchSonic: (a) Segmentation of user study data into 4.5-minute-long episodes, (b) Detection of food intake within each episode.
5.2.1. Eating Episode Detection
To evaluate MunchSonic in detecting eating episodes and intake counts within episodes, we segmented the 45-minute data collected from each participant in the user study into 10 segments of equal length (4.5 minutes each). According to the definition provided in (Thomaz et al., 2015a; Zhang et al., 2022), if a segment of length $t_{w}$ contains $\frac{t_{w}+20}{5}$ intakes, which is 5 intakes for the aforementioned segments, then that segment is defined as an eating episode. To evaluate MunchSonic’s efficacy in detecting these eating episodes, we deployed a majority voting strategy where a segment is labeled as an eating episode if 50% of the frames are labeled as either food_intake or chewing. The number of undetected eating episodes for the MunchSonic dataset is 1 out of 43 ground truth episodes, resulting in a False Negative Rate of 0.023. Additionally, there were 77 non-eating episodes in the dataset, of which 2 were incorrectly detected as eating by MunchSonic, leading to a False Positive Rate of 0.026.
5.2.2. Food Intake Counting
Furthermore, to count the number of intakes in each segment, we increment the intake count for that segment by 1 if we find one food_intake frame followed by at least two chewing frames or windows within the next 3.0 seconds. We compute the Mean Absolute Error (MAE) between the ground truth number of intakes and the predicted count as a metric to evaluate this. The average ground truth intake count in one 4.5-minute eating segment (containing more than 5 ground truth intakes within the timeframe) and non-eating segment (containing fewer than 5 ground truth intakes) is 17.65 and 1.19, respectively. The MAE for counting intakes in the eating segments is 2.01, leading to an average error of 11.39% in predicting the number of eating intakes. For the non-eating segments, the MAE for counting intakes is 0.047, leading to a mean error of 3.95% in predicting the intake counts during non-eating episodes.
5.2.3. Chewing Time Estimation
To evaluate the coverage of predicted chewing time, we calculate the Mean Absolute Error (MAE) between the total ground truth chewing time in a 4.5-minute episode and the predicted chewing time, measured in seconds, across all 12 user study participants. The average chewing time across eating episodes is 163.6 seconds (2.73 minutes) and 9.17 seconds for non-eating episodes in the MunchSonic dataset. Interpolating from the frame-level predictions of MunchSonic, the MAE of the estimated chewing time for eating episodes is 12.87 seconds, leading to a coverage of 92.13% of chewing time. For the non-eating episodes, the MAE of the estimated chewing time is 1.61 seconds, leading to a coverage of 82.5%. This higher coverage across non-eating episodes indicates MunchSonic’s potential in detecting snacking events, which was outside the scope of the user study.
6. Discussion
6.1. Analyzing the Passive Sensed Ultrasonic Range for Tracking Dietary Actions
Figure 6. Differential echo profiles of actions with active and passive acoustic sensing. The $x$-axis represents time, and the $y$-axis represents the echo distance from the MunchSonic form factor, ranging from 0 cm to 50 cm.
Theoretically, our system focuses on nearly inaudible sound frequencies above 18 KHz, which should not interfere with most daily activities, including eating. However, some sounds generated by eating can potentially have frequency components above 18 KHz. To evaluate the impact of these passively sensed ultrasonic signals on distinguishing dietary actions, we conducted a preliminary study with one researcher as a participant. In this study, we turned off both speakers of MunchSonic to prevent them from transmitting any C-FMCW chirps and performed the same set of activities as in the MunchSonic user study. The microphones on both sides continued to record the surrounding acoustic signals while the researcher conducted similar dietary actions. We then created differential echo profile windows (illustrated in the bottom row of Figure 6) in the same way as before and trained the MunchSonic deep learning model. The mean cross-session macro F1-score for this user-dependent model was 0.322. In contrast, when we collected similar data from the same participant with the speakers transmitting ultrasonic chirps, the mean cross-session macro F1-score was 0.976. As illustrated in Figure 6, and evident from the performance difference mentioned, active acoustic sensing plays a crucial role in distinguishing dietary activities by tracking movements of body parts (e.g., face, arms) involved in dietary actions. While passive ultrasonic signals might have some minor impact on detecting fine-grained events, quantifying this impact numerically is challenging and beyond the scope of this paper. Our experiments indicate that chewing crunchy foods (e.g., chips) generated more ultrasonic components in the spectrogram. However, the signal strength of reflections from the ultrasonic chirps transmitted by MunchSonic is much higher than the passively sensed components, potentially overshadowing their feature importance in MunchSonic’s deep learning model training. Further investigation into this will be left for future work.
6.2. Impact of Sensing Range and Temporal Context
Figure 7. Ablation study to measure the impact of sensing range and sliding window size on the performance of MunchSonic. Here, all the sliding window sizes mentioned have a 50% overlap.
We evaluate the impact of the sensing range and temporal context for the MunchSonic system in this experiment. Figure 7 demonstrates that a sliding window size of 2.00 seconds with a 50% overlap yields the best tracking performance. We also observe that a sensing range of 50 to 80 cm demonstrates optimal performance in tracking fine-grained dietary actions. This can be attributed to the fact that most movements related to the dietary actions MunchSonic tracks happen in the jaw and mouth region, which is within 30 cm of the eyeglasses form factor. Additionally, MunchSonic tracks hand movements for food intake, so extending the sensing range to 50 cm yields the best performance. The poorer performance at longer ranges can be rationalized by the fact that various activities can occur in that region, which may not be related to eating.
6.3. Potential Application Scenarios
We envision MunchSonic being potentially useful in various contexts, including eating behavior assessment, chronic disease management, and overall well-being. Our system can enhance eating behavior assessment by providing a passive and objective measurement of eating episodes and micro-level actions (e.g., eating pauses, drinking, chewing) in real-world environments. By accurately classifying these actions, we can derive the ”microstructure of meals” (Bellisle, 2020), including eating duration, eating speed, chewing rate, and chewing efficacy, and their variations across episodes. Traditionally, these metrics are assessed through semi-structured interviews (e.g., Eating Disorder Examination (Fairburn and Wilson, 1993; for Research on Eating Disorders at Oxford, CREDO)) or self-report questionnaires (e.g., (Sasaki et al., 2003)), which rely on recall and approximation and are prone to biases. Our system offers an automated, precise, and naturalistic approach to capturing these behaviors. Clinical studies indicate that these metrics are linked to various health outcomes. For instance, higher eating and chewing rates are associated with higher body mass index (e.g., obesity (Sonoda et al., 2018)), eating disorders (Walsh et al., 1989), and metabolic diseases (Yuan et al., 2021). Therefore, precise assessment of these metrics could enhance the monitoring and management of these health outcomes over time, facilitating real-time monitoring and interventions.
7. Conclusion
MunchSonic demonstrates accurate tracking of fine-grained dietary actions in real-world settings, surpassing other systems with its single instrumentation point, high precision, privacy preservation, and scalability. Our work holds significant potential for advancing dietary health assessment and monitoring in both technical and healthcare domains.